OpenStudy (anonymous):

If the diameter of the circle is 6 cm, find its circumference. [Use 3.14 for π.]

Nnesha (nnesha):

\[\huge\rm C = \pi \times d\]when they gave you diameter you can use this equation :-)

OpenStudy (anonymous):

@Nnesha wow ilysm

Nnesha (nnesha):

hmmm huh ?? *gO_Ogle*

Nnesha (nnesha):

ohhh thanks ;D :- ) :-)

OpenStudy (anonymous):

@Nnesha haha your welcome

Nnesha (nnesha):

:-) let me know what you get :-)

OpenStudy (anonymous):

@Nnesha 18.84?

Nnesha (nnesha):

yep!! :)

OpenStudy (anonymous):

@nnesha thank youu, but i have another question . could you naswer it?

Nnesha (nnesha):

btw \(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\)

OpenStudy (anonymous):

thank youuu lol

Nnesha (nnesha):

nope i'm not gonna answer it i'll try my best to explain and then u have to find answer :- )

OpenStudy (anonymous):

Find the supplement of an angle that measures 20 degrees.

OpenStudy (anonymous):

got chaa lol

Nnesha (nnesha):

supplementary angle add up to what degree !!! ????

OpenStudy (anonymous):

uhhhh, 140?

OpenStudy (anonymous):

90?

OpenStudy (anonymous):

160?

Nnesha (nnesha):

oppos which one :D ?? btw all that wrong :( :( so all supplementary that add up to 180

Nnesha (nnesha):

so if one angle measure 20 degree you have to subtract 20 nto 180 180 -20= ???

OpenStudy (anonymous):

160? :D

Nnesha (nnesha):

yep!!!

OpenStudy (anonymous):

Thank youuu :)

Nnesha (nnesha):

my pleasure!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! :-) _-_ o^_^o gO_Od luck!!